OpenStudy (anonymous):

In the first quadrant representative rectangle so that has a vertex at the origin and the opposite vertex on the parabola y = x ^ 2 +3. Determine the dimensions of the rectangle so that its area is maximized.

OpenStudy (anonymous):

@hartnn @hba

hartnn (hartnn):

 so the rectangle has dimension of \(x \times (x^2+3)\) so, your area \(A =x \times (x^2+3)=x^3+3x\) can you maximize this ? differentiate it and set A'=0 to get values of 'x'